HW7 010 Homework • Unanswered Price Quantity Refer to HW7 Q9. A firm in a perfectly competitive market can sell any quantityHW7 Q11 Homework • Unanswered Price Quantity A monopoly firm has to drop the price of its product if it wants to sell more quHW7 Q12 Homework • Unanswered Price Quantity A monopoly firm has to drop the price of its product if it wants to sell more quAnswer only, No explanation needed
Answer #1

a 100 P = $.5 0:=4 Of = 6 2. Marginal Revenue = A Total Revenue ! a polf-plli - Pl0lf-Qi) I . Margina Renenue = 516-4) =$10 I0-12 Y = ч Фі ч ; (- 3 Ф 6 marginal Revenue - A Total Revenue – С. Ф. - f; © •rjin few,че – 36) - (42t) = ч 1